An object's height, in meters, when thrown vertically from the ground at a velocity of 80m/s is given by the equation H=-4.9t^2+80t. Ignoring wind resistance, what is the height of the ball after 2 seconds? 3 seconds? 10 seconds?
In this problem I substituted the seconds in the given equation, but I was told that was incorrect....HELP PLEASE!!!

Answer by stanbon(75887)

An object's height, in meters, when thrown vertically from the ground at a velocity of 80m/s is given by the equation H=-4.9t^2+80t. Ignoring wind resistance, what is the height of the ball after 2 seconds? 3 seconds? 10 seconds?
In this problem I substituted the seconds in the given equation, but I was told that was incorrect....HELP PLEASE!!!
-----
2 seconds
H(2) = -4.9(2^2) + 80(2) = 140.4 meters
-------
Do the same for 3 seconds and for 10 seconds.
=============
cheers,
Stan H.